# Multi-task Representation Learning for Mixed Integer Linear Programming

### v1

@@ 1 Introduction
Many real-world problem domains, such as path planning [46], scheduling [17], and network design [10, 25], fall into the category of combinatorial optimization (CO) and are generally NP-hard to solve. Designing efficient algorithms for CO problems is both important and challenging. Mixed Integer Linear Programs (MILPs) provide a versatile framework for modeling and solving various CO problems. MILPs involve optimizing a linear objective function subject to linear constraints, with some variables restricted to integer values. Significant advancements in MILP solvers, such as Gurobi [20] and SCIP [5], have been achieved by leveraging techniques like Branch-and-Bound (BnB) [36], complemented by a suite of heuristics to enhance performance.
@@ 2 Background
This section defines MILPs and provides background knowledge on our three tasks: Configuration, Backdoor, and PaS.
@@ 3.1 Machine Learning for MILP Solving
There has been extensive research leveraging machine learning techniques to improve MILP solvers. Common approaches represent MILPs as bipartite graphs [19] and employ graph neural networks (GCN [19, 12, 45, 32, 52, 27, 24] and GAT [26, 7]) to learn various MILP decisions. Learning methods are diverse, some commonly used are imitation learning [22, 49, 45], contrastive learning [26, 7, 27, 39], and reinforcement learning [50, 48]. For a comprehensive survey on machine learning for MILP solving, we refer readers to [47]. Here, we focus on our selected three tasks.
@@
Solution prediction: The goal is to predict partial assignments of high-quality feasible solutions in a MILP to guide the search. [12] identifies variables that remain unchanged across near-optimal solutions and searches within their neighborhood. [43] and [32] propose fixing predicted variables and letting the MILP solver optimize the rest as a warm start. However, if predictions are inaccurate, fixing variables can lead to low-quality or infeasible solutions. [21] introduces PaS, which searches for solutions within a predefined neighborhood of the prediction, improving feasibility and quality. [27] extends PaS using contrastive learning and novel optimization-based methods for handling low-quality or infeasible solutions.
@@ 3.2 ML-guided solving Generalization
Despite advancements in ML-guided solving, generalizing learned models across tasks and problem domains remains a key challenge. Most ML methods are tailored to specific problem classes, limiting their applicability in real-world scenarios. Recently, some works have addressed this issue. [28] introduced Distributional MIPLIB, the first multi-domain library for advancing ML-guided MILP methods and exploring cross-domain generalization. [38] proposed MILP-Evolve, leveraging large language models to generate diverse MILP classes, showing strong generalization when trained on a large dataset. [14] developed GOAL, a generalist model capable of solving various problems in CO like routing, scheduling, packing, and graph problems, and fine-tuning to new problem domains. Additionally, multi-task learning has been applied in other CS problems like computer vision [41] and natural language processing [8], but no one has applied it to the CO domain.
@@ 4 Multi-task Representation Learning
The pipeline is shown in Figure 1. With this multi-task training pipeline, we achieve a standard representation embedding of MILPs that can be generalized across different tasks within the same problem domain. It is worth noting that this pipeline is adaptable to any graph representation of MILPs, any size-invariant network architecture, and any loss function. This paper focuses on the bipartite graph representation, Graph Attention Networks, and contrastive loss. In the following subsections, we explain the data collection process, bipartite graph representation, network architecture, loss function, and applying learned networks for the tasks: Backdoor, PaS, and Configuration.
@@ 5.1 Experiment Setup
We utilize Distributional-MIPLIB [28] to generate 200 training S instances and 100 test S instances for each task. We also generated another 100 test L instances for each task to test the model’s generalizability. We train and fine-tune the model on S instances and test on both S and L instances. Each task involves different problem sizes: Backdoor focuses on improving optimal solving time for smaller instances (with optimal solutions found in approximately hundreds of seconds), while PaS and Configuration address more challenging instances (solving optimally takes over an hour), aiming to achieve better primal solutions within a runtime cutoff. Table 1 provides detailed information on the generated instances for each problem domain and task.
@@
Table 2: Solve time (seconds) for Backdoor and Primal Integral for PaS averaged over 100 test instances for each benchmark. We compare the performance of Gurobi, Single-task (trained on Backdoor or PaS), and Multi-task-BAPAS. Results include the mean, standard deviation, and the number of instances each approach wins. The best-performing entries are highlighted in bold for clarity.
@@
(a) CA-S (2000, 4000)
(b) MIS-S (5, 6000)
(c) MVC-S (5, 6000)
(d) CA-L (3000, 6000)
(e) MIS-L (5, 9000)
(f) MVC-L (5, 9000)
Figure 2: The Primal Gap (the lower, the better) as a runtime function averaged over 100 test instances on PaS and each benchmark. We compare the performance of Gurobi (green line), Single-task on PaS (blue line), and Multi-task-BAPAS (orange line).
@@
(a) CA-S (2000, 4000)
(b) MIS-S (4, 3000)
(c) MVC-S (4, 3000)
(d) CA-L (3000, 6000)
(e) MIS-L (5, 6000)
(f) MVC-L (5, 6000)
Figure 3: New task performance: The Primal Gap (the lower, the better) as a runtime function averaged over 100 test instances on Configuration and each benchmark. We compare the performance of SCIP (green line), SMAC (red line), Single-task on Configuration (blue line), and Multi-task-BAPAS (orange line).
@@ 5.2 Results
Table 3: New task performance: Solve time (seconds) for Backdoor and Primal Integral for PaS and Configuration averaged over 100 test instances for each benchmark. We compare the performance of Gurobi, Single-task, and correspond Multi-task. Results include the mean and the number of instances each approach wins. The number of Wins does not add up to 100 for Configuration because SMAC wins in some instances. The reforming entries are highlighted in bold for clarity.
@@
New Task Performance: We compare SCIP, SMAC, Single-task, and Multi-task-BAPAS on Configuration using S and L instances from CA, MIS and MVC. Results are presented in Figure 3 and the table with the primal integral is available in the Appendix. First, in our new task of learning to generate configurations, Single-task outperforms in the primal gap over baselines on CA-S and MVC-S, and also achieving 18.27% and 66.97% improvements in the primal integral over SMAC. While it initially generates worse solutions than the baselines on MIS-S, it surpasses them by the end of the runtime cutoff. This demonstrates that our new task can effectively learn to generate configurations that enhance solving performance. Now focusing on Multi-task-BAPAS, we fine-tune only the task-specific layer using the same 200 configuration instances as the Single-task, while keeping the rest of the network architecture trained on Backdoor and PaS. Surprisingly, the multi-task model performs similarly to, or even better than, the Single-task on S instances. Multi-task-BAPAS achieves better solution progress within the first 600 seconds on CA-S and MVC-S and consistently outperforms the Single-task throughout the runtime on MIS-S. For primal integral, Multi-task-BAPAS reaches an average of 11.3% over Single-task across three S benchmarks.
@@
To further validate the robustness of our approach, we performed cross-task evaluations to compare the learned MILP embeddings against Single-task. Specifically, we evaluated Multi-task-PASCONFIG on Backdoor, Multi-task-PASCONFIG on PaS, and Multi-task-BAPAS on Configuration, assessing their generalization performance on 100 large instances from each benchmark. Table 3 highlights consistent improvements of the multi-task models over the solver and Single-task across all benchmarks. These experiments demonstrate that the success of our multi-task training framework is not limited to a single pair of tasks but is broadly effective. Additional tables and figures for S and L instances and comparisons among different multi-task models on the same task are included in the Appendix.
@@ Acknowledgments
The National Science Foundation (NSF) supported the research under grant number 2112533: "NSF Artificial Intelligence Research Institute for Advances in Optimization (AI4OPT)".

### v2

@@ 1 Introduction
Many real-world problem domains, such as path planning [48], scheduling [19, 9], and network design [12, 27], fall into the category of combinatorial optimization (CO) and are generally NP-hard to solve. Designing efficient algorithms for CO problems is both important and challenging. Mixed Integer Linear Programs (MILPs) provide a versatile framework for modeling and solving various CO problems. MILPs involve optimizing a linear objective function subject to linear constraints, with some variables restricted to integer values. Significant advancements in MILP solvers, such as Gurobi [22] and SCIP [5], have been achieved by leveraging techniques like Branch-and-Bound (BnB) [38], complemented by a suite of heuristics to enhance performance.
@@ 2 Background
This section defines MILPs and provides background knowledge on our three tasks: Backdoor, PaS, and Configuration.
@@ 3.1 Machine Learning for MILP Solving
There has been extensive research leveraging machine learning techniques to improve MILP solvers. Common approaches represent MILPs as bipartite graphs [21] and employ graph neural networks (GCN [21, 14, 47, 34, 55, 29, 26] and GAT [28, 7]) to learn various MILP decisions. Learning methods are diverse, some commonly used are imitation learning [24, 51, 47], contrastive learning [28, 7, 29, 41], and reinforcement learning [52, 50, 8]. For a comprehensive survey on machine learning for MILP solving, we refer readers to [49]. Here, we focus on our selected three tasks.
@@
Solution prediction: The goal is to predict partial assignments of high-quality feasible solutions in a MILP to guide the search. [14, 54] identifies variables that remain unchanged across near-optimal solutions and searches within their neighborhood. [45] and [34] propose fixing predicted variables and letting the MILP solver optimize the rest as a warm start. However, if predictions are inaccurate, fixing variables can lead to low-quality or infeasible solutions. [23] introduces PaS, which searches for solutions within a predefined neighborhood of the prediction, improving feasibility and quality. [29] extends PaS using contrastive learning and novel optimization-based methods for handling low-quality or infeasible solutions.
@@ 3.2 ML-guided solving Generalization
Despite advancements in ML-guided solving, generalizing learned models across tasks and problem domains remains a key challenge. Most ML methods are tailored to specific problem classes, limiting their applicability in real-world scenarios. Recently, some works have addressed this issue. [30] introduced Distributional MIPLIB, the first multi-domain library for advancing ML-guided MILP methods and exploring cross-domain generalization. [40] proposed MILP-Evolve, leveraging large language models to generate diverse MILP classes, showing strong generalization when trained on a large dataset. [16] introduced GOAL, a generalist model for various combinatorial optimization problems, but it has not been applied to MILP domains. Additionally, multi-task learning has been applied in other CS problems like computer vision [43] and natural language processing [10], but no one has applied it to the CO domain.
@@ 4 Multi-task Representation Learning
The pipeline is shown in Figure 1. With this multi-task training pipeline, we achieve a standard representation embedding of MILPs that can be generalized across different tasks within the same problem domain. It is worth noting that this pipeline is adaptable to any graph representation of MILPs, any size-invariant network architecture, and any loss function. This paper focuses on the bipartite graph representation, Graph Attention Networks, and contrastive loss. In the following subsections, we explain the data collection process, bipartite graph representation, network architecture, loss function, and applying learned networks for the tasks: Backdoor, PaS, and Configuration. We choose these three MILP tasks because there are existing works [7, 29] that have utilized the same architecture and loss function, which allows us to effectively explore a multitask model without starting entirely from scratch. Additionally, these three tasks require only one-time inference with static features.
@@ 5.1 Experiment Setup
We utilize Distributional-MIPLIB [30] to generate 200 training small (S) instances and 100 test S instances for each task. We also generated another 100 test large (L) instances for each task to test the model’s generalizability. We train and fine-tune the model on S instances and test on both S and L instances. Each task involves different problem sizes: Backdoor focuses on improving optimal solving time for smaller instances (with optimal solutions found in approximately hundreds of seconds), while PaS and Configuration address more challenging instances (solving optimally takes over an hour), aiming to achieve better primal solutions within a runtime cutoff. Table 1 provides detailed information on the generated instances for each problem domain and task.
@@
Table 2: Same-Task Performance: Solve time (seconds) for Backdoor and Primal Integral for PaS averaged over 100 test instances for each benchmark. We compare the performance of Gurobi, Single-task (trained on Backdoor or PaS), and Multi-task-BAPAS. Results include the mean, standard deviation, and the number of instances each approach wins. The best-performing entries are highlighted in bold for clarity.
@@
Figure 2: Same-Task Performance: The Primal Gap (the lower, the better) as a runtime function averaged over 100 test instances on PaS and each benchmark. We compare the performance of Gurobi (green line), Single-task on PaS (blue line), and Multi-task-BAPAS (orange line).
@@
Figure 3: New-task Performance: The Primal Gap (the lower, the better) as a runtime function averaged over 100 test instances on Configuration and each benchmark. We compare the performance of SCIP (green line), SMAC (red line), Single-task on Configuration (blue line), and Multi-task-BAPAS (orange line).
@@ 5.2 Results
Table 3: Cross-task Evaluations: Solve time (seconds) for Backdoor and Primal Integral for PaS and Configuration averaged over 100 test instances for each benchmark. We compare the performance of Gurobi, Single-task, and correspond Multi-task. Results include the mean and the number of instances each approach wins. The number of Wins does not add up to 100 for Configuration because SMAC wins in some instances. The reforming entries are highlighted in bold for clarity.
@@
New-Task Performance:
We compare SCIP, SMAC, Single-task, and Multi-task-BAPAS on Configuration using S and L instances from CA, MIS and MVC. Results are presented in Figure 3 and the table with the primal integral is available in the Appendix. First, in our new task of learning to generate configurations, Single-task outperforms in the primal gap over baselines on CA-S and MVC-S, and also achieving 18.27% and 66.97% improvements in the primal integral over SMAC. While it initially generates worse solutions than the baselines on MIS-S, it surpasses them by the end of the runtime cutoff. This demonstrates that our new task can effectively learn to generate configurations that enhance solving performance. Now focusing on Multi-task-BAPAS, we fine-tune only the task-specific layer using the same 200 configuration instances as the Single-task, while keeping the rest of the network architecture trained on Backdoor and PaS. Surprisingly, the multi-task model performs similarly to, or even better than, the Single-task on S instances. Multi-task-BAPAS achieves better solution progress within the first 600 seconds on CA-S and MVC-S and consistently outperforms the Single-task throughout the runtime on MIS-S. For primal integral, Multi-task-BAPAS reaches an average of 11.3% over Single-task across three S benchmarks.
@@
To further validate the robustness of our approach, we performed cross-task evaluations on the learned MILP embeddings against Single-task. Specifically, we evaluated Multi-task-PASCONFIG on Backdoor, Multi-task-BACONFIG on PaS, and Multi-task-BAPAS on Configuration, assessing their generalization performance on 100 large instances from each benchmark. Table 3 highlights consistent improvements of the multi-task models over the solver and Single-task across all benchmarks. These experiments demonstrate that the success of our multi-task training framework is not limited to a single pair of tasks but is broadly effective. Additional tables and figures for S and L instances and comparisons among different multi-task models on the same task are included in the Appendix.
@@ Acknowledgments
We would like to thank the anonymous reviewers for their constructive feedback to improve this paper. The National Science Foundation (NSF) supported the research under grant number 2112533: "NSF Artificial Intelligence Research Institute for Advances in Optimization (AI4OPT)".